OpenStudy (adll23):

A system of equations is shown below: x + y = 3 2x - y = 6 The x-coordinate of the solution to this system of equations is _____. (4 points)

jhonyy9 (jhonyy9):

@adll23 do you wan collaborate please ?

OpenStudy (adll23):

i think so @bibby

OpenStudy (adll23):

?? @jhonyy9

jhonyy9 (jhonyy9):

ok how you that is the right easy way ? than you see this system

OpenStudy (raffle_snaffle):

You can solve this one of 3 ways.

jhonyy9 (jhonyy9):

how you think

OpenStudy (adll23):

what are the 3 ways @raffle_snaffle

OpenStudy (adll23):

idk @jhonyy9

OpenStudy (raffle_snaffle):

You can either add them or you can solve using system of equations.

OpenStudy (adll23):

oh i see ill add them

OpenStudy (raffle_snaffle):

show me you work please=)

OpenStudy (adll23):

1x+1y=3 2x-1y=6 it equals 3x+0y=9 is this right?

OpenStudy (adll23):

@raffle_snaffle

OpenStudy (jdoe0001):

hmm if you solve the 2nd one, 2x-y=6, for say... "y", what does it give you anyway?

OpenStudy (raffle_snaffle):

yep

OpenStudy (adll23):

idk i am really bad at math

OpenStudy (raffle_snaffle):

Okay lets solve using the other method.

OpenStudy (raffle_snaffle):

lets take x+y = 3. you can solve for x or y which one would you like to solve for?

OpenStudy (adll23):

x

OpenStudy (adll23):

@raffle_snaffle

jhonyy9 (jhonyy9):

@adll23 please write your calculi here

OpenStudy (adll23):

idk how

jhonyy9 (jhonyy9):

how you have got these results by elimination method

jhonyy9 (jhonyy9):

x+y=3 2x-y=6

OpenStudy (raffle_snaffle):

subtract y from both sides.